# Aggregated Attributions for Explanatory Analysis of 3D Segmentation Models

### v2

@@ Acknowledgments
This work was financially supported by the Polish National Center for Research and Development grant number INFOSTRATEG-I/0022/2021-00. We thank Mateusz Krzyzinski and Paulina Tomaszewska for valuable feedback on the initial version of this work.
@@ B.3 Implementation details of baseline model for 3D Image Segmentation
We use a cosine annealing learning rate scheduler [43]. We implemented our network in Python v3.9 using PyTorch v2.1 and the MONAI library [9]. As an evaluation metric, we use the Dice Similarity Coefficient. We employ a one-way analysis of variance to evaluate the significance of variations among the segmentation performance metrics. We use $p<0.05$ to distinguish a statistically significant difference.
@@ C.1 Analysis of outliers with Agg2Exp
Table 2: P-values and correlation coefficients from Spearman correlation test between Dice and anomaly scores of the IF models.
@@ Appendix D Comparison of the segmentation performance between the baseline models
Table 3: TotalSegmentator test set performance comparison of mean Dice Score Coefficient (mDSC) for the following classes. For better readability, we present the left and right lungs as means of their subparts and autochthon as one class: Aorta (A), Left Lung (LL), Right Lung (RL), Trachea (T), Heart (H), Pulmonary Vein (PV), Thyroid Gland (TG), Ribs (R), Vertebrae (V), Autochthon (AU), Sternum (S), Costal Cartilages (CC). *indicates statistical significance between Swin UNETRv2 and other state-of-the-art methods mDSC ($p<0.05$). **indicates training with self-supervised pre-trained weights.

### v3

@@ Acknowledgments
This work was financially supported by the Polish National Center for Research and Development grant number INFOSTRATEG-I/0022/2021-00.
@@ B.3 Implementation details of baseline model for 3D Image Segmentation
We use a cosine annealing learning rate scheduler [44]. We implemented our network in Python v3.9 using PyTorch v2.1 and the MONAI library [10]. As an evaluation metric, we use the Dice Similarity Coefficient (DSC). We employ a one-way analysis of variance to evaluate the significance of variations among the segmentation performance metrics. We use $p<0.05$ to distinguish a statistically significant difference.
@@ C.1 Analysis of outliers with Agg2Exp
Table 2: P-values and correlation coefficients from Spearman correlation test between DSC and anomaly scores of the IF models.
@@ Appendix D Comparison of the segmentation performance between the baseline models
Table 3: TotalSegmentator test set performance comparison of mean DSC for the following classes. For better readability, we present the left and right lungs as mean of their subparts and autochthon as one class: Aorta (A), Left Lung (LL), Right Lung (RL), Trachea (T), Heart (H), Pulmonary Vein (PV), Thyroid Gland (TG), Ribs (R), Vertebrae (V), Autochthon (AU), Sternum (S), Costal Cartilages (CC). *indicates statistical significance between Swin UNETRv2 and other state-of-the-art methods mDSC ($p<0.05$). **indicates training with self-supervised pre-trained weights.